OpenStudy (anonymous):

8(6-5x)=-24(x-2)-16x

OpenStudy (a_clan):

48 - 40x = -24x +48 -16x x=0

OpenStudy (anonymous):

48-40x=-24x+48-16x (expand brackets) 48x-40x=48-40x Hence the statement is true.

jhonyy9 (jhonyy9):

48-40x=-24x +48 -16x 48-40x = 48-40x